# Reducing Task Discrepancy of Text Encoders for Zero-Shot Composed Image Retrieval

### v1

@@ 3.2 Target-anchored text contrastive learning
Figure 2 illustrates the overview of our RTD. We use CLIP backbone and pre-trained projection module $\phi$ produced by the existing projection-based ZS-CIR methods [15, 16, 17]. The text encoder is trained using the proposed loss function (Eq. 1) while applying the refined batch sampling and concatenation scheme. During inference, the procedure mirrors that of existing ZS-CIR methods, except we utilize the updated text encoder $\psi_{T}^{tr}$ instead of the frozen one $\psi_{T}$. Note that our method only updates the text encoder while the image encoder and the projection module are frozen.
Remark. We note that our entire process, including text triplet construction and the training step, is efficient due to the advantages of language-only training as highlighted in [17]. First, for the text triplet construction, we only require an efficient text triplet generation process without the resource-intensive text-to-image generation phase [8, 12]. If we choose the template-based text triplet generation process, it becomes even more efficient by eliminating the need for the LLM generation step. Moreover, our generated text triplets occupy just 100MB, whereas storing a similar quantity of images requires significantly more space (e.g., around 400GB in the case of CC3M [6]). Second, the training complexity for the text encoder is substantially lower than that for the visual encoder due to the relatively short token lengths of texts ($\sim$12) compared to images (256). Consequently, the average inference time of the CLIP ViT-L/14 image encoder is $\times$ 3.5 times slower than that of the text encoder. In Section 4.4, we demonstrate a more efficient implementation option by selectively updating only a few layers of the text encoder. Namely, we verify that the size of the learnable parameters of the text encoder can be reduced to the same size as the parameters of the projection module $\phi$ while achieving similar results.
3.3 Can RTD really reduce the task discrepancy of the text encoder?
Table 1: T2I retrieval performance of different text encoders on CIRCO validation dataset.
@@ 3.3 Can RTD really reduce the task discrepancy of the text encoder?
In this subsection, we quantitatively verify whether RTD really can reduce the task discrepancy. We first conduct a toy experiment that measures the text-to-image (T2I) retrieval performance of the text encoder with the modification instruction. We retrieve the target images $I_{t}$ with the concatenated text query $T_{r+c}$ or the ideal target caption $T_{t}$. If our text encoder successfully handles the discrepancy due to the concatenated caption, the text encoder updated by RTD will perform better than the frozen one. We use the CLIP ViT-L/14 and CIRCO [16] validation dataset for evaluation. Since the CIRCO dataset only has CIR triplets $(I_{r},T_{c},I_{t})$, we use the BLIP [22] captioner to generate $T_{r}$ and $T_{t}$ corresponding to the $I_{r}$ and $I_{t}$, respectively. Here, the simple concatenation scheme is applied for the text query $T_{r+c}$ in all cases for a fair comparison. Table 1 shows that when the text encoder is frozen, the retrieval results using the concatenated caption $T_{r+c}$ are significantly worse than those using the target caption $T_{t}$. It supports the claim that the frozen text encoder suffers from the negative effects of task discrepancy between the pretext task and the CIR task. In contrast, the text encoder updated by RTD shows a significant improvement over the frozen text encoder, showing that it successfully reduces the task discrepancy.

### v2

@@ 3.2 Target-anchored text contrastive learning
Fig. 2 illustrates the overview of RTD. We use CLIP backbone and pre-trained projection module $\phi$ produced by the existing projection-based CIR methods. The text encoder is trained using the proposed loss function (Sec. 3.2) while applying the refined batch sampling and concatenation scheme. During inference, the procedure mirrors that of existing projection-based CIR methods, except we utilize the updated text encoder $\psi_{T}^{tr}$ instead of the frozen one $\psi_{T}$. Note that our method only updates the text encoder while the image encoder and the projection module are frozen.
Remark 1: Low data acquisition cost. The text triplets we use can be obtained at a significantly lower cost than CIR triplets. First, the text triplet generation process requires only an easily obtainable caption dataset for generation, whereas other approaches that generate CIR triplets [45, 22, 48] necessitate image or video datasets along with an additional collection phase for semantically similar images or videos. Second, text triplet generation can avoid resource-intensive text-to-image generation [16, 5], making it 15$\times$ faster than the CIR triplet generation process used in CompoDiff [16]. Furthermore, if we choose the rule-based text triplet generation strategy, the process becomes 570$\times$ faster than the full CIR triplet generation with images; 1M text triplets generation takes just 0.1 hours. Lastly, the storage for the generated text triplets takes up only 100MB, whereas storing a similar quantity of images requires significantly more space (e.g., 400GB for CC3M [39]). More details can be found in App. C.
Table 1: Training and inference time comparisons. Purple denotes the additional resource required for RTD. Inference time is measured on a single A100 GPU for a single image. Note that the training hours for SEARLE, LinCIR for 8 A100 GPUs [17] and their RTD variants are trained on 4 A100 GPUs. CompoDiff is trained on 128 A100 and MagicLens is trained on 128 TPUs.
Method CLIP Backbone Training (h) Inference (s)
SEARLE L/14 4.2h 0.02s
SEARLE (+ RTD) L/14 4.58h (+0.38h) 0.02s
LinCIR L/14 0.5h 0.02s
LinCIR (+ RTD) L/14 0.88h (+0.38h) 0.02s
LinCIR G/14 0.8h 0.05s
LinCIR (+ RTD) G/14 3.3h (+2.5h) 0.05s
CompoDiff L/14 231h 0.12s
MagicLens L/14 6h unknown
CIReVL L/14 - 3.02s
Remark 2: Training and inference efficiency. As reported in Tab. 1, the additional training cost of RTD is small—0.38 hours for ViT-L/14 on 4 A100 GPUs. In contrast, CompoDiff and MagicLens require 231 hours on 128 A100 GPUs and 6 hours on 128 TPUs, respectively, making the cost of RTD negligible. Even with a larger ViT-G/14 backbone, training RTD takes only 2.5 hours on 4 A100 GPUs—still faster than synthetic CIR triplet-based methods and even comparable to SEARLE in ViT-L/14 (4.2h). This training efficiency mainly stems from the advantage of text-only training. As highlighted by Gu et al. [17], the training complexity for the text encoder is remarkably lower than that for the visual encoder due to the relatively short token lengths of texts ($\approx$12) compared to images (256). The average inference time of the CLIP ViT-L/14 image encoder is $\times$ 3.5 times slower than that of the text encoder. Another source of efficiency is that RTD requires relatively few iterations (approximately 2000), as the text encoders are already pre-trained and only require minor adjustments to learn the relationships between text triplets. Moreover, in Sec. B.4, we present a more efficient implementation by selectively updating only a few layers of the text encoder. Along with efficient training, RTD retains the fast inference speed of the original projection-based methods, remaining 150$\times$ faster than training-free CIReVL.
4 Experiments
4.1 Experimental setup
Implementation details. We use the AdamW optimizer [29] with a weight decay of 0.01. The learning rate is set to $10^{-5}$, with a batch size of 512. We select the text encoder model with the best zero-shot CIRR [28] dev R@1 score for evaluating RTD. We mainly use the visual and textual encoders of the CLIP ViT-L/14 [34] as our backbone. Unless otherwise noted, we use LLM-based 2.5M text triplets provided by CompoDiff [16] for the training. We set the $\tau$ as $0.07$ in Sec. 3.2 and scale the standard deviation of Gaussian distribution as $0.5$ for the noise injection. More results on various noise distributions can be found in the Sec. B.7. All experiments were conducted using four NVIDIA A100 GPUs with Python 3.8 and Pytorch [32].
Evaluation datasets and metrics. We compare projection-based CIR methods on five benchmark datasets: CIRR [28], CIRCO [4], FashionIQ [46] (FIQ), COCO object composition [36], and GeneCIS [44]. Details of each dataset are in the Sec. A.1. For CIRR, FIQ, COCO, and GeneCIS, we have reported their recall scores at the top K retrieval results (R@K). Since the CIRCO dataset includes multiple positive images for each query, we use a ranking-based metric—mean Average Precision scores at the top K results (mAP@K) [30, 10]. For the main results, we compare the results on the FIQ validation split, as well as the test sets of CIRR and CIRCO. For the ablation studies and analyses, the validation splits of these three datasets are utilized. GeneCIS and COCO object composition results and their detailed explanations can be found in the Sec. B.1.
Baselines. We evaluate the effect of our method when combined with publicly available projection-based ZS-CIR methods: Pic2Word [36], SEARLE [4], LinCIR [17], Context-I2W [42], and FTI4CIR [26]. All these methods share the similar core concept shown in Fig. 1, but use different training schemes. Pic2Word[36] optimizes contrastive loss between the image embedding and its projected text embedding of “a photo of [\$]” to obtain the projection module $\phi$. Similarly, SEARLE [4] employs a two-stage approach, starting with an optimization-based textual inversion phase followed by a distillation phase for the projection module $\phi$. LinCIR [17] introduces a language-only self-supervised task involving keyword token replacement by letting the original text embedding and the replaced text embedding whose keyword tokens are changed to the projected original text embedding by $\phi$. Context-I2W [42] refine the projection module by selecting relevant visual information [42]. Lastly, FTI4CIR [26] separately maps images into subject- and attribute-oriented pseudo-word tokens [26]. Details on how RTD is combined with them can be found in Sec. A.3.
Table 2: Comparison with other baselines. Note that this comparison is not entirely fair due to differences in backbone models and training data. Purple denotes the performance gain from our method, while red and blue highlight the best and second-best scores, respectively.
Method Backbone CIRR CIRCO FashionIQ
R@1 R@5 R@10 mAP@5 mAP@10 mAP@25 R@10 R@50
Pic2Word [36] CLIP ViT-L/14 $24.2$ $51.5$ $64.1$ $8.3$ $9.1$ $10.1$ 25.3 44.9
+RTD $27.9$ ($\mathbf{+3.6}$) $56.2$ ($\mathbf{+4.8}$) $68.5$ ($\mathbf{+4.4}$) $9.1$ ($\mathbf{+0.9}$) $9.6$ ($\mathbf{+0.5}$) $10.7$ ($\mathbf{+0.6}$) 27.6 ($\mathbf{+2.3}$) 48.9 ($\mathbf{+4.0}$)
SEARLE [4] CLIP ViT-L/14 $24.9$ $52.3$ $65.7$ $11.6$ $12.7$ $14.3$ $25.0$ $45.3$
+RTD $26.6$ ($\mathbf{+1.7}$) $56.2$ ($\mathbf{+3.9}$) $69.0$ ($\mathbf{+3.3}$) $16.5$ ($\mathbf{+4.9}$) $17.9$ ($\mathbf{+5.2}$) $19.8$ ($\mathbf{+5.4}$) $29.3$ ($\mathbf{+4.4}$) $50.7$ ($\mathbf{+5.4}$)
Context-I2W [42] CLIP ViT-L/14 25.6 55.4 68.6 - - - 27.9 49.1
+RTD $29.2$ ($\mathbf{+4.6}$) $58.4$ ($\mathbf{+3.0}$) $70.5$ ($\mathbf{+1.9}$) - - - $28.1$ ($\mathbf{+0.2}$) $49.5$ ($\mathbf{+0.4}$)
FTI4CIR [26] CLIP ViT-L/14 25.9 55.6 67.7 15.1 16.3 18.1 29.4 50.9
+RTD $29.1$ ($\mathbf{+3.2}$) $58.8$ ($\mathbf{+3.2}$) $70.4$ ($\mathbf{+2.7}$) $16.2$ ($\mathbf{+1.1}$) 17.4 ($\mathbf{+1.1}$) 19.4 ($\mathbf{+1.3}$) $30.6$ ($\mathbf{+1.2}$) $51.7$ ($\mathbf{+0.8}$)
LinCIR [17] CLIP ViT-L/14 $23.8$ $52.9$ $66.5$ $13.0$ $14.1$ $15.8$ $27.4$ $47.7$
+RTD $26.6$ ($\mathbf{+2.9}$) $56.2$ ($\mathbf{+3.3}$) $69.0$ ($\mathbf{+2.5}$) $17.1$ ($\mathbf{+4.1}$) $18.1$ ($\mathbf{+4.0}$) $20.1$ ($\mathbf{+4.3}$) $30.2$ ($\mathbf{+2.8}$) $51.1$ ($\mathbf{+3.4}$)
LinCIR [17] FSC-CLIP [31] ViT-L/14 33.6 63.1 74.5 14.2 15.4 17.0 28.3 49.5
+RTD $35.9$ ($\mathbf{+2.3}$) $\mathbf{67.5}$ ($\mathbf{+4.4}$) $78.1$ ($\mathbf{+3.6}$) $18.3$ ($\mathbf{+4.1}$) $19.7$ ($\mathbf{+4.3}$) $21.3$ ($\mathbf{+4.3}$) $31.2$ ($\mathbf{+2.9}$) $52.3$ ($\mathbf{+2.8}$)
KEDs [41] CLIP ViT-L/14 26.4 54.8 67.2 - - - 26.8 47.9
CIReVL [20] CLIP ViT-L/14 24.5 52.3 64.9 18.6 19.0 20.9 28.6 48.6
LDRE [47] CLIP ViT-L/14 26.5 55.6 67.5 23.4 24.0 26.4 28.5 50.5
MT-CIR [7] CLIP ViT-L/14 25.5 54.6 67.6 10.4 11.6 13.0 35.4 57.4
MagicLens [48] CLIP ViT-L/14 30.1 61.7 74.4 29.6 30.8 33.4 30.7 52.5
Compodiff [16] CLIP ViT-L/14 26.7 55.0 72.6 12.6 13.4 15.8 36.0 48.6
CoVR [45] BLIP ViT-L/14 39.3 68.2 78.9 - - - 27.7 44.6
CASE [22] BLIP ViT-L/14 35.4 65.8 78.5 - - - - -
LinCIR [17] CLIP ViT-G/14 34.9 64.5 76.1 20.6 21.9 24.1 44.5 65.5
+RTD 36.3 ($\mathbf{+1.4}$) 67.5 ($\mathbf{+3.0}$) 78.3 ($\mathbf{+2.2}$) 21.1 ($\mathbf{+0.5}$) 22.3 ($\mathbf{+0.4}$) 24.5 ($\mathbf{+0.4}$) 46.2 ($\mathbf{+1.7}$) 67.3 ($\mathbf{+1.8}$)
Compodiff [20] CLIP ViT-G/14 34.7 64.3 75.1 15.3 17.7 19.4 39.0 51.7
CIReVL [20] CLIP ViT-G/14 26.7 55.1 74.5 26.8 27.6 30.0 32.2 52.4
LDRE [47] CLIP ViT-G/14 36.2 66.4 77.3 31.1 32.2 35.0 32.5 55.4
We train these methods with the CLIP ViT-L/14 in our main experiments. To further verify the compatibility of RTD, we also report results for the ViT-B/32 backbone with SEARLE, Pic2Word, and LinCIR. Moreover, we conduct experiments with a larger backbone (ViT-G/14) and a fine-tuned CLIP model for enhanced compositionality (FSC-CLIP [31]) for LinCIR, enabled by its fast training capability. We use the publicly available pre-trained model for SEARLE (ViT-B/32, ViT-L/14), Pic2Word (ViT-L/14), Context-I2W (ViT-L/14), and FTI4CIR (ViT-L/14). Otherwise, we reproduce the results using the official implementation. When reproducing, we adhere to the same settings in the original papers. For example, we select the final last epoch model for the Pic2Word ViT-B/32 model and choose the model based on the best zero-shot CIRR dev R@1 score for LinCIR. We additionally compare our method with a diverse set of CIR approaches, including another projection-based ZS-CIR method (KEDs [41]), another attempt to address task discrepancy (MT-CIR [7]), approaches leveraging synthetically generated CIR triplets (CoVR [45], CASE [22], MagicLens [48] and Compodiff [16]), and the training-free methods (CIReVL [20], LDRE [47]).
4.2 Main results
Tab. 2 shows the overview of comparison results with state-of-the-art CIR methods. First, we assess the impact of integrating RTD with existing projection-based CIR methods (SEARLE, Pic2Word, LinCIR, Context-I2W, and FTI4CIR) across CIRR, CIRCO, and FashionIQ. RTD consistently boosts performance, yielding an average improvement of over 2.78 points. This trend also holds for the GeneCIS and COCO object composition datasets, as detailed in Sec. B.1.
Second, we compare the integration of RTD with those leveraging synthetically generated CIR triplets (CoVR, CASE, MagicLens, and Compodiff). We focus on this comparison to ensure fairness, as leveraging text triplets and updating the textual backbone of RTD may not be fully aligned with the standard projection-based CIR setting. We observe that RTD delivers competitive or superior performance compared to these resource-intensive synthetic CIR triplet-based methods, while being significantly more efficient—over 10–100$\times$ faster, as noted in Remark-2 of Sec. 3. For example, with the same CLIP ViT-L/14 backbone, LinCIR + RTD outperforms MagicLens and CoVR in FashionIQ R@10 and R@50 while FTI4CIR + RTD or Context-I2W + RTD achieve performance on par with or exceeding that of MagicLens, MT-CIR, and Compodiff in CIRR metrics. When combined with larger (ViT-G/14) or fine-tuned (FSC-CLIP) backbones, RTD achieves the best or second-best result on the CIRR and FashionIQ benchmarks. We believe this flexibility of using different backbones underscores the advantage of efficient text-only training of RTD, inherited by LinCIR [17]. We reemphasize that even with a larger backbone like ViT-G/14, RTD remains significantly faster in training than existing synthetic triplet-based methods, as highlighted in Tab. 1.
Lastly, across both backbones (ViT-L/14, G/14), LinCIR + RTD mostly outperforms training-free CIReVL and LDRE, except on CIRCO metrics. As shown in Tab. 1, CIReVL is significantly slower (150$\times$) than projection-based methods (including RTD), while LDRE incurs even greater inference time due to its ensemble-based strategy.
Table 3: Ablation study. Unlike in Tab. 2, for ablation studies and analyses, validation splits of three CIR datasets are used for evaluation. We measure the impact of TCL loss (Sec. 3.2), refined batch sampling (RB), and refined concatenation scheme (RC). All models are based on LinCIR ViT-L/14. The first row denotes the vanilla LinCIR without RTD. “Avg” denotes the average of all reported metrics. Bold indicates the best result.
TCL RB RC CIRR (R) CIRCO (mAP) FIQ (R) Avg
Text Anc @5 @10 @10 @25 @10 @50
- - 54.3 67.8 12.7 14.5 27.4 47.7 37.4
$(T_{r},T_{r})$ 56.0 69.7 13.4 15.2 28.2 48.8 38.5
$(T_{r+c},T_{t})$ 58.2 71.5 14.4 16.0 26.9 47.9 39.2
$(T_{r+c},T_{t})$ 58.2 71.3 15.0 16.7 27.4 49.3 39.6
$(T_{r+c},T_{t})$ 54.3 67.0 12.2 13.6 25.0 45.3 36.3
$(T_{r+c},T_{t})$ 57.9 71.1 16.1 17.8 30.2 51.1 40.7
4.3 Ablation studies
Tab. 3 presents the effectiveness of the proposed components: target-anchored text contrastive loss (TCL), refined batch sampling (RB), and refined concatenation scheme (RC). All evaluation results are on the validation splits. All model variants use ViT-L/14 and a projection module $\phi$ from LinCIR, making the results in row 1 indicative of the original performance of LinCIR. We first compare the impact of the text pairs fed into TCL loss. We compare our design choice $(T_{r+c},T_{t})$ (from the generated text triplets) with $(T_{r},T_{r})$, which is the sole option for constructing a pair given a single conventional caption $T_{r}$. The results demonstrate that, on average, using generated triplets (3rd row) is more effective than using original conventional text pairs (2nd row), particularly in the CIRR and CIRCO datasets. In addition, RB (4th row) and RC (6th row) significantly enhance the overall performance, demonstrating the effectiveness of these components. Finally, we measure the impact of using the frozen text encoder for target caption $T_{t}$, denoted as “Anchor” in the table. Significant performance degradation is observed when the learnable text encoder is used for extracting the embedding of the target caption $T_{t}$ (5th row) compared to the target-anchored case (4th row), supporting the importance of the anchoring design choice.
Table 4: T2I retrieval performance of different text encoders on CIRCO val set. “Update (pair)” refers to the setting in the second row of Tab. 3, which uses contrastive learning with the pair $(T_{r},T_{r})$. Bold indicates the best result, excluding the first row (oracle case).
@@ 4.3 Ablation studies
4.4 Anaylses on our core motivation
We conduct experiments to validate our core motivation (reducing task discrepancy) and verify that the observed gains stem from it. Other details follow Sec. 4.3.
@@ 4.4 Anaylses on our core motivation
Can RTD really reduce the task discrepancy of the text encoder? We first quantitatively verify whether RTD indeed reduces the task discrepancy. We first conduct a controlled experiment that measures the text-to-image (T2I) retrieval performance of the text encoder with conditional texts. We retrieve the target images $I_{t}$ with the concatenated text query $T_{r+c}$ or the ideal target caption $T_{t}$. If our text encoder successfully handles the discrepancy due to the concatenated caption, the text encoder updated by RTD will perform better than the frozen one or “Update (pair)”, which is updated by contrastive learning with the $(T_{r},T_{r})$ pair (corresponding to row 2 of Tab. 3). We use the CLIP ViT-L/14 and CIRCO [4] validation dataset for evaluation. Since the CIRCO dataset only has CIR triplets $(I_{r},T_{c},I_{t})$, we use the BLIP [23] captioner to generate $T_{r}$ and $T_{t}$ corresponding to the $I_{r}$ and $I_{t}$, respectively. Here, the simple concatenation scheme is applied for the text query $T_{r+c}$ in all cases for a fair comparison. Tab. 4 shows that when the text encoder is either frozen or updated with $(T_{r},T_{r})$, the retrieval results using the concatenated caption $T_{r+c}$ are significantly worse than those using the target caption $T_{t}$. It supports the claim that the frozen text encoder suffers from the negative effects of task discrepancy between the pretext and CIR tasks. Moreover, it suggests that simply updating the text encoder with $(T_{r},T_{r})$ pair fails to reduce this discrepancy. In contrast, the text encoder updated by RTD shows a significant improvement over the frozen text encoder or one updated with $(T_{r},T_{r})$, showing that it successfully reduces the task discrepancy.